Please help me solve this problem, can you please show and explain the work? I feel I am missing a very simple step. Thank you
What is the future value of $3000 in a bank account for 9 years at 6% compounded bimonthly? Round to the nearest cent.

Answer by rfer(16322)

Every other month or twice a month?
-----------------------------------
twice a month
FV=3000(1+0.06/24)^24*9
FV=3000(1.0025)^216
FV=3000(1.71485)
FV=$5144.55
-----------------
every other month
FV=3000(1+0.06/6)^6*9
FV=3000(1.01)^54
FV=3000(1.71141)
FV=$5134.23